OpenStudy (anonymous):

Chemical analysis of a germanium crystal reveals indium at a level of 0.0003091 atomic percent. Assuming that the concentration of thermally excited charge carriers from the Ge matrix is negligible, what is the density of free charge carriers (free carriers/cm3) in this Ge crystal?

OpenStudy (aaronq):

"Each In atom will attract an electron and thus create a “mobile hole”; we only have to determine the number of In atoms/cm3. The atomic volume of the host crystal (Ge) is given on your PT as 13.57 cm3/mole. # Ge atoms/cm3 = 6.02 ×1023 atoms × 1 mole 1 mole 13.57 cm3 = 4.44 × 1022 atoms/cm3 # In atoms/cm3 = 4.44 × 1022 × 0.0003 × 10-2 = 1.33 × 1017 In/cm3 The number of free charge carriers (“holes”) is 1.33 × 1017/cm3; they are created through the acquisition of one electron by each In atom from the valence band of the host crystal." http://web.mit.edu/3.091/www2/Fall10/homework/homework6_with_solutions_2010.pdf i didn't write this but i was interested so i looked it up.